# 3D sans 3D Scans: Scalable Pre-training from Video-Generated Point Clouds

### v1

@@ Appendix C Further Analysis Experiments
kNN neighborhood size (see Table LABEL:tab:knn). We analyze the effect of the neighborhood size used to construct the kNN graph for the Laplacian smoothing loss. In our LAM3C, the Laplacian smoothing loss is computed on a kNN graph constructed for each video-generated point cloud. The value of $k$ determines the size of the local neighborhood, thus controlling the strength and spatial extent of the smoothing imposed by the regularization.
We pre-train LAM3C with three values $k{=}\{8,24,32\}$, and evaluate the resulting models on ScanNet semantic segmentation. As shown in Table LABEL:tab:knn, k=24k{=}24 yields the best performance, particularly in the Linear Probing setting where representation quality is most directly reflected. We consider that smaller $k$ results unstable local neighborhoods, while larger $k$ leads to over-smoothing.
Radius threshold for noisy point removal (see Table LABEL:tab:radius). We analyze the effect of the radius threshold used to remove distant neighbors when constructing the kNN graph for the Laplacian smoothing loss. In video-generated point clouds, a noisy reconstructed point may cause a query point to retrieve far-away points among its top-k neighbors, which violates the locality assumption required for Laplacian regularization. To mitigate this issue, we apply a distance threshold $r_{max}$ and discard neighbors whose Euclidean distance exceeds this value.
@@
Semantic Seg. Test/Val set ScanNet [dai2017scannet] ScanNet200 [rozenberszki2022language] ScanNet++ Val [yeshwanth2023scannet++] S3DIS Area 5 [Armeni_2016_CVPR]
@@
Huber loss parameter $\delta$ (see Table LABEL:tab:huber_loss). We analyze the effect of the Huber loss parameter $\delta$, which determines the transition point between the quadratic and linear regimes of the Huber loss. In our formulation, the Laplacian smoothing loss penalizes the difference between a point feature and the mean of its neighbors. When this residual is small, the Huber loss applies an $L_{2}$ penalty; when it exceeds $\delta$, it switches to an $L_{1}$ penalty to reduce the influence of outliers.
@@
Gaussian scale parameter $\sigma$ (see Table LABEL:tab:sigma). We analyze the effect of the Gaussian scale parameter $\sigma$, which controls the decay rate of the distance-based weighting used in the Laplacian smoothing loss. We assign each edge a distance-based weight $w_{ij}=\exp(-{\|p_{i}-p_{j}\|^{2}}/{\sigma^{2}})$, where a larger $\sigma$ results in a slower decay of the weights, allowing distant points to retain non-negligible influence. Conversely, a smaller $\sigma$ produces a sharper decay, restricting the effective neighborhood to only very close points. To account for the varying density of video-generated point clouds, our method adaptively sets $\sigma$ as the median of the kNN distances for each point cloud.
@@
Loss weights $\lambda$ and $\mu$ (see Table LABEL:tab:loss_weight). We analyze the impact of the weighting coefficients $\lambda$ and $\mu$ used for the Laplacian smoothing loss and Noise consistency loss, respectively. As defined in Eq. (5), the overall objective consists of the distillation loss from Sonata and two additional regularization terms. Here, we focus on the balance between these two regularized losses, which are introduced to stabilize representation learning from video-generated point clouds. In our formulation, the coefficient $\lambda$ for the Laplacian smoothing loss follows a scheduled progression during training, whereas the coefficient $\mu$ for the Noise consistency loss is fixed.
@@ Appendix D Fairness considerations for baseline
. This setup is not comparable to our LAM3C, which, even during finetuning evaluations, does not see any {test / val} data. Such {test / val} inclusion has been common practice in previous indoor 3D-SSL methods [wu2023masked, wu2024towards], but blurs the true generalisation performance measurement.

### v2

@@ Appendix C Further Analysis Experiments
kNN neighborhood size (see Table LABEL:tab:knn). We analyze the effect of the neighborhood size used to construct the kNN graph for the Laplacian smoothing loss. In our LAM3C, the Laplacian smoothing loss is computed on a kNN graph constructed for each VGPC. The value of $k$ determines the size of the local neighborhood, thus controlling the strength and spatial extent of the smoothing imposed by the regularization.
We pre-train LAM3C with three values $k{=}\{8,24,32\}$, and evaluate the resulting models on ScanNet semantic segmentation. As shown in Table LABEL:tab:knn, k=24k{=}24 yields the best performance, particularly in the Linear Probing setting where representation quality is most directly reflected. We consider that smaller $k$ results in unstable local neighborhoods, while larger $k$ leads to over-smoothing.
Radius threshold for noisy point removal (see Table LABEL:tab:radius). We analyze the effect of the radius threshold used to remove distant neighbors when constructing the kNN graph for the Laplacian smoothing loss. In VGPC, a noisy reconstructed point may cause a query point to retrieve far-away points among its top-k neighbors, which violates the locality assumption required for Laplacian regularization. To mitigate this issue, we apply a distance threshold $r_{max}$ and discard neighbors whose Euclidean distance exceeds this value.
@@
Semantic Seg. Test/Val set ScanNet [15] ScanNet200 [35] ScanNet++ Val [56] S3DIS Area 5 [2]
@@
Huber loss parameter $\delta$ (see Table LABEL:tab:huber_loss). We analyze the effect of the Huber loss parameter $\delta$, which determines the transition point between the quadratic and linear regimes of the Huber loss. In our formulation, the Laplacian smoothing loss penalizes pairwise feature differences between neighboring points on the kNN graph. When the feature difference on an edge is small, the Huber penalty behaves quadratically; when it exceeds $\delta$, it becomes linear, reducing the influence of outlier edges.
@@
Gaussian scale parameter $\sigma$ (see Table LABEL:tab:sigma). We analyze the effect of the Gaussian scale parameter $\sigma$, which controls the decay rate of the distance-based weighting used in the Laplacian smoothing loss. We assign each edge a distance-based weight $w_{ij}=\exp(-{\|p_{i}-p_{j}\|^{2}}/{\sigma^{2}})$, where a larger $\sigma$ results in a slower decay of the weights, allowing distant points to retain non-negligible influence. Conversely, a smaller $\sigma$ produces a sharper decay, restricting the effective neighborhood to only very close points. To account for the varying density of VGPC, our method adaptively sets $\sigma$ as the median of the kNN distances for each point cloud.
@@
Loss weights $\lambda$ and $\mu$ (see Table LABEL:tab:loss_weight). We analyze the impact of the weighting coefficients $\lambda$ and $\mu$ used for the Laplacian smoothing loss and Noise consistency loss, respectively. As defined in Eq. (5), the overall objective consists of the distillation loss from Sonata and two additional regularization terms. Here, we focus on the balance between these two regularized losses, which are introduced to stabilize representation learning from VGPC. In our formulation, the coefficient $\lambda$ for the Laplacian smoothing loss follows a scheduled progression during training, whereas the coefficient $\mu$ for the Noise consistency loss is fixed.
@@
[…]
Figure 9: Component synergy on ScanNet semantic segmentation (linear probing). Replacing VGGT with $pi^{3}$, adding indoor alignment, and introducing LAM3C each improve performance, with further gains from scaling RoomTours and using PTv3-Large.
@@ Appendix D Fairness considerations for baseline
. This setup is not comparable to our LAM3C, which, even during finetuning evaluations, does not see any {test / val} data. Such {test / val} inclusion has been common practice in previous indoor 3D-SSL methods [53, 52], but blurs the true generalisation performance measurement.
@@
Appendix E Discussion
Our findings show that high-fidelity 3D scans are not a prerequisite for effective 3D-SSL. Despite being noisy, incomplete, and inconsistent, VGPC reconstructed from unlabeled videos prove sufficient for 3D-SSL. Perhaps most notably, LAM3C performs strongly in linear probing, a setting that has not been widely explored in previous 3D-SSL methods. This indicates that these imperfect reconstructions still encode rich geometric cues relevant to real-world tasks. Overall, our results highlight a promising direction toward scalable 3D data generation. Instead of costly real 3D scans, unlabeled videos offer a far more scalable source of 3D supervision, where the key is learning to handle noisy and missing regions effectively.
Appendix F Limitations and outlook
Although the reconstruction models used in this paper use real point clouds during training, our core contribution lies in showing that previously untapped resources such as web videos can be used for 3D-SSL and become more useful at scale. By piggybacking on reconstruction models, our approach benefits from the fast progress in this domain and might enable unifying 3D reconstruction and recognition models. Our method is designed for predominantly static indoor scenes and may be limited in highly dynamic environments, where reconstruction quality degrades due to many moving objects. While our noise regularization mitigates some outliers in RoomTours, explicit modeling of dense dynamics and motion remains an important direction for future work.